# High-performance real-world optical computing trained by in situ gradient-based model-free optimization

### v5

@@ Abstract
Computational Optics; Optical Neural Network; Model-free Training;
@@ 1 Introduction
  • Application scenario: We demonstrate, as a proof-of-concept, that our G-MFO-trained optical computing system’s effectiveness on classifying four types of white blood cells from their phase maps in a marker-free111 Herein, we use ”marker-free” but not ”label-free” to avoid confusion. This is because ”label-free” in biomedical research and high-level vision have different meanings. manner with a testing accuracy of $73.8\%$, making it a promising approach for image-free, marker-free, and high-speed cell analysis (Sec. 4.4).
Open source. Upon acceptance, we will open-source all the code (including simulators, algorithm-related, hardware-related, and baselines) related to our work to facilitate research on training optical computing systems. We hope to promote the reproducibility and follow-ups for this hardware-centric research direction.
@@
TABLE I: Comparison of strategies on training optical computing systems along the axes of in situ training capability (in situ), in silico computation overhead (computation overhead), the requirement on the recording of the input object to the task (input image free), required in silico training time, the dimensionality of trainable parameters (dimensionality), requirements on intermediate measurement and whether or not use gradient-based update that is more efficient (gradient). *Denotes compatibility with adding nonlinearity to the optical computing system. Simulator-based training (SBT), Hybrid model-based training (H-MBO), Learned model-based training (L-MBO), Intermediate measurement optical backpropagation (IMOB), Forward-forward training (FFT), Genetic algorithm (GA), Gradient-based model-free optimization (G-MFO).
@@ 2.1 In situ training strategies to optimize the optical computing system
The majority of research into optical computing system training has historically relied on in silico training. This process involves carrying out forward and backward propagation calculations on an external computer that simulates the physical system through a digital twin [9, 12]. However, this technique can lead to discrepancies between the simulation and actual reality due to inaccuracies in the physical system’s representation.
@@
In recent years, progress has been made in developing in situ training strategies that use data gathered from real-world optical computing systems to mitigate the reality gap and improve experimental performance. As a first attempt, we categorize previous and our in-situ training methods into two genres: model-free optimization (MBO) and model-based optimization (MFO) methods. We differentiate these two genres by whether they use numerical models of the optical computing system during the parameter updating process. A tabular comparison of the previous and our methods is in Tab. I. Details of the illustrations and comparisons are illustrated below.
@@
MBO methods. Physics-aware training (PAT) [15], hybrid training (HBT) [16] or adaptive training (AT) [20] (H-MBO) are one type of the model-based methods that conduct forward pass in a real system and back-propagate the gradients with the aid of a physics-based simulator. Backpropagation through learned model (L-MBO) is another method that gathers the real data from the optical computing system to train a neural proxy of the optical computing system [21, 22]. With such a learned proxy rather than the physics-based model utilized in PAT/HBT/AT, the backward parameter update process can reduce the bias from the model mismatch between the simulator and the real system. During training, one has to optimize both the task’s and proxy’s parameters to optimize the task performance and realize high fidelity. The task performance for the MBO, especially the L-MBO, is high, while back-propagating with the digital twin also burdens the training process. Moreover, all the MBO approaches require images of input objects during the training, imposing another burden on the overall process.
@@
MFO methods. Unlike the aforementioned MBO methods, MFO methods do not require constructing a digital twin of the optical computing system; they directly use the real system’s output or intermediate measurements to update its parameters. Population-based algorithms such as genetic algorithm (GA) methods [30] use the output from the optical computing system to optimize parameters by simulating the process of natural selection. Still, the genetic algorithm’s performance scales poorly to the dimensionality of parameters [31]. Forward-forward training (FFT) is a greedy multi-layer learning procedure for mortal computation [32] and it has been applied to train optical computing systems reently [26, 27]. However, labels must be added to each input sample during the training and testing. As a result, such training strategies require masking the input, which might be unrealistic in many scenarios, and the testing process of forward-forward training is prolonged. Intermediate measurement optical backpropagation (IMOB) proposed using the adjoint variable method or direct feedback alignment, etc., to calculate gradients [23, 24, 25]. However, such methods require intermediate system measurements for every layer’s output, which might be infeasible or require additional measurement setups.
@@
We propose an MFO training method using Monte Carlo gradient estimation to automatically update the optical computing system parameters. Our work shares similarities with Zhang et al., which uses the genetic algorithm for weight update [30], as our and their methods both generate a batch of sampled parameters during the training process. The difference is that our parameter updating process is more efficient owing to a score gradient estimator rather than their heuristic hill-climbing type algorithm. Moreover, our method does not need specific conditions, such as changing the system or manipulating the input object, which are more or less required in other MBO and MFO methods.
@@ 3.3 In situ G-MFO
$=\frac{1}{M}\sum_{j=1}^{M}[\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(f_{sys}(x_{i},w_{j}),y_{i})]\nabla_{\theta}\log{{p(w_{j}|\theta)}}.$ (8b)
Here we define $r(w_{j})=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(f_{sys}(x_{i},w_{j}),y_{i})$ as the negative reward corresponding to each weight $w_{j}$. We use Equation 8b as the score gradient estimator for G-MFO, where the score function is $\nabla_{\theta}\log{p(w_{j}|\theta)}$, which has been widely used in other areas, such as policy gradient algorithms in reinforcement learning [19] and diffusion models [43].
@@
Figure 3: An illustration of the training process for the MFO experiment on the real system.
Figure 4: Setups of optical computing systems. (a) In the single-layer optical computing system, we have a light source incident from the top left, passing through the collimator lens (CL), polarizer (P), and the first beam splitter (BS1) and then reflecting from the input layer, which displays phase object by phase delay from the SLM1. The reflected light then passes through BS1 and beam splitter 2 (BS2) and arrives at the optical computing layer, where we use SLM2 to upload the optical processing weights. The reflected light from the optical computing layer then arrives at the output layer, and we calculate the task-specific loss based on the arrival signal. We label the last two paths as $Prop_{IC}$ and $Prop_{CO}$, respectively, in the figure. (b) In the two-layer optical computing system, the light passing through CL and P is reflected by a mirror (M) and directed towards the input layer. The input layer displays binary intensity objects using a DMD. The reflected light then reaches the optical computing layer 1 and, subsequently, optical computing layer 2, where we use SLM1 and SLM2 to upload the optical processing weights. Finally, the light arrives at the output layer, and we obtain the output of the optical computing system. We label the three paths as $Prop_{IC_{1}}$, $Prop_{C_{1}C_{2}}$, $Prop_{C_{2}O}$.
3.4 Experiment and simulation detail of our diffractive optical computing system
3.4.1 Experimental setups of the real single-layer and two-layer optical computing systems $f_{sys-1-layer}$ and $f_{sys-2-layer}$.
We build a single-layer optical computing system and a two-layer optical computing system to validate the effectiveness of the proposed G-MFO-based training strategy (see Fig. 4 and Supplement Fig. S.1).
Single-layer optical computing system: A laser field $u_{laser}$ propagates onto the input layer (SLM1), where the phase-valued object $u_{obj}$ is displayed. The light field reflects from the input layer onto the optical computing layer $u_{w}$ (SLM2). A camera then detects the diffraction light at the output layer. The length $d_{IC}$ and $d_{CO}$ for $Prop_{IC}$ and $Prop_{CO}$ in Fig. 4(a) are $215.1mm$ and $201.6mm$, respectively. As there are multiple components in the system (e.g., SLMs and cameras), it is difficult to manually align them with high accuracy. We developed a digital alignment method based on the homography technique [45] using SLMs to compensate for the alignment errors when we implement the baseline method HBT on the real single-layer optical computing system; otherwise, the performance will be pretty poor. Details of our digital alignment method are in Supplement Sec. S.5.
Double-layer optical computing system: A DMD displays the binary intensity object $u_{obj}$ on the input layer. The light field is then reflected from the input layer onto the optical computing layer1 $u_{w1}$ (SLM1) and subsequently to the optical computing layer2 $u_{w2}$ (SLM2). Following this, a camera captures the modulated light present at the output layer. The length $d_{C_{1}C_{2}}$ and $d_{C_{2}O}$ for $Prop_{C_{1}C_{2}}$ and $Prop_{C_{2}O}$ in Fig. 4(b) are $215.1mm$ and $201.6mm$, respectively. More details of the optical computing system are in Supplement Sec. S.1.
3.4.2 Differentiable physics-based simulators $\hat{f}_{sys-1-layer}$ and $\hat{f}_{sys-2-layer}$.
${\color[rgb]{1,0,0}\hat{f}_{prop}}(u_{in},z):u_{out}=\mathcal{F}^{-1}(\mathcal{F}(u_{in})\times\mathcal{F}(h_{prop}(z))),$ (10a)
${\color[rgb]{0.390625,0.515625,0.9609375}\hat{f}_{mod}}(u_{in},u_{element}):u_{out}=u_{in}*u_{element},$ (10b)
${\color[rgb]{0.984375,0.6328125,0.234375}\hat{f}_{det}}(u_{in}):I_{out}=|u_{in}|^{2},$ (10c)
where $h_{prop}(z)=\frac{e^{jkz^{\prime}}}{j\lambda z^{\prime}}e^{\frac{jk}{2z}(x^{\prime 2}+y^{\prime 2})}$ is the propagation kernel under Fresnel approximation [46] with a propagation distance $z^{\prime}$, wavelength $\lambda$ and angular wave number $k$, $u_{element}$ denotes the wavefront modulation from the programmable optical devices, $\mathcal{F}$ denotes the Fourier transform.
@@ 3.4 Experiment and simulation detail of our diffractive optical computing system
Figure 5: Visualization of the optical computing system’s experimental outputs and confusion matrices trained with G-MFO. (a) An input phase object digit ’$2$’ from the MNIST dataset is modulated by the optical computing layer with weight $w$ trained using G-MFO. The system correctly predicts the input as digit ‘$2$’, as the output image has the largest intensity at the region corresponding to digit ’$2$’. (b) Confusion matrix on the MNIST dataset with a training accuracy of $83.1\%$. (c) An example of a ’pullover’ from the FMNIST dataset is correctly predicted. (d) Confusion matrix on the FMNIST dataset with a training accuracy of $74.0\%$.
Simulator of a single-layer optical computing system. Based on the modules in Eq. 10, the simulator of a single-layer $\hat{f}_{sys-1-layer}=\{{\color[rgb]{0.390625,0.515625,0.9609375}\hat{f}_{mod}},{\color[rgb]{1,0,0}\hat{f}_{prop}},{\color[rgb]{0.390625,0.515625,0.9609375}\hat{f}_{mod}},{\color[rgb]{1,0,0}\hat{f}_{prop}},{\color[rgb]{0.984375,0.6328125,0.234375}\hat{f}_{det}}\}$ is constructed by chaining the building blocks ${1-5}$:
$1.u_{out}={\color[rgb]{0.390625,0.515625,0.9609375}\hat{f}_{mod}}(u_{laser},u_{obj}),$ (11a)
$2.u_{out}={\color[rgb]{1,0,0}\hat{f}_{prop}}(u_{out},d_{IC}),$ (11b)
$3.u_{out}={\color[rgb]{0.390625,0.515625,0.9609375}\hat{f}_{mod}}(u_{out},u_{w}),$ (11c)
$4.u_{out}={\color[rgb]{1,0,0}\hat{f}_{prop}}(u_{out},d_{CO}),$ (11d)
$5.I_{cam}={\color[rgb]{0.984375,0.6328125,0.234375}\hat{f}_{det}}(u_{out}).$ (11e)
@@ 4 Results
  • HBT: The hybrid training strategies with the digitally aligned model. A detailed description of our implementation of the HBT method is in Supplement Sec. S.6.
Training details of all simulations and experiments are in Supplement Sec. S.3.
@@ 4.1 General performance evaluation on the MNIST and FMNIST dataset
The simulation outcomes for the MNIST and FMNIST datasets are in Table 2. The ideal testing accuracy reaches $89.5\%$ on MNIST and $80.5\%$ on FMNIST. However, the presence of misalignment reduces the accuracy to $78.4\%$ and $59.8\%$ using HBT, a decrease of approximately $10\%$ and $20\%$. In contrast, the G-MFO method hits an accuracy of $87.0\%$ and $74.1\%$, effectively mitigating the detrimental impact of system misalignment.
@@
TABLE II: Numerical performance comparison on MNIST and FMNIST datasets on two-layer diffractive optical computing system.
Tab. III quantitatively shows that our method achieves higher classification accuracy than the HBT and SBT methods on both datasets in the experiments. The SBT method performs poorly due to the reality gap between the simulator and the real system. The HBT method suffers from the bias between $\hat{f}_{sys}$ and ${f}_{sys}$ in the backward process, while the G-MFO bypasses the bias-sensitive modeling and updates gradients solely with $f_{sys}$. We further show G-MFO’s capability to fine-tune the HBT result. The last row of Tab. III shows that the unbiased G-MFO method further improves the results of HBT. Moreover, we also empirically find that fine-tuning outperforms G-MFO only. Fig. 5 visualizes some experimental outputs and confusion matrices using the G-MFO method.
@@
TABLE III: Experimental performance comparison on MNIST and FMNIST datasets on the single-layer optical computing system. Results of the ideal mode are from the simulator, whose parameters are determined from experiments while we impose no misalignment on the simulator. The lower four rows are experimental results from SBT, HBT, G-MFO, and HBT+G-MFO (G-MFO fine-tuning upon the result from HBT). Our method outperforms the SBT and HBT methods on the MNIST and FMNIST datasets in experiments.
@@
We also perform experiments on training a real two-layer optical computing system $f_{sys-2-layer}$ on the MNIST dataset. For this experiment, we utilize a dataset comprising 10,000 instances for training, another 10,000 for validation, and a separate set of 10,000 for testing. The training, validation and testing accuracy are $80.43\%$, $81.1\%$, and $80.3\%$, respectively.
When compared with the simulated G-MFO outcomes presented in Table. II, the experimental accuracy is lower. We hypothesize that this discrepancy is attributed, in part, to mechanical disturbances encountered during the multi-day span of the experiment. Moreover, the training accuracy from two-layer system is lower than that of single-layer system while the validation and testing accuracy is higher. This is because we use more samples when training two-layer system, which alleviates overfitting.
@@
4.2 Advantage: memory- and computation-efficient training enabled by G-MFO.

### v6

@@ Abstract
Computational Optics; Optical Neural Network; Model-free Optimization;
@@ 1 Introduction
  • Application scenario: We demonstrate, as a proof-of-concept, that our G-MFO-trained optical computing system’s effectiveness on classifying four types of white blood cells from their phase maps in a marker-free111 Herein, we use ”marker-free” but not ”label-free” to avoid confusion. This is because ”label-free” has different meanings in biomedical research and high-level vision. manner with a testing accuracy of $73.8\%$, making it a promising approach for image-free, marker-free, and high-speed cell analysis (Sec. 4.4).
Open source. Our code will be available at repository Model-free-Computational-Optics to facilitate further research and applications on computational optics and training of optical computing systems. We hope to promote the reproducibility and follow-ups for this hardware-centric research direction.
@@
TABLE I: Comparison of strategies on training optical computing systems along the axes of in situ training capability (in situ), in silico computation overhead (computation overhead), the requirement on the recording of the input object to the task (input image free), required in silico training time, the dimensionality of trainable parameters (dimensionality), requirements on intermediate measurement and whether or not use gradient-based update that is more efficient (gradient). Simulator-based training (SBT), Hybrid model-based training (H-MBO), Learned model-based training (L-MBO), Intermediate measurement optical backpropagation (IMOB), Forward-forward training (FFT), Trial and error, Genetic algorithm (GA), Gradient-based model-free optimization (G-MFO).
@@ 2.1 In situ training strategies to optimize the optical computing system
Most research into optical computing system training has historically relied on in silico training. This process involves forward and backward propagation calculations on an external computer that simulates the physical system through a digital twin [9, 12]. However, this technique can lead to discrepancies between the simulation and actual reality due to inaccuracies in the physical system’s representation.
@@
In recent years, progress has been made in developing in situ training strategies that use data gathered from real-world optical computing systems to mitigate the reality gap and improve experimental performance. As a first attempt, we categorize previous and our in-situ training methods into two genres: model-based optimization (MBO) and model-free optimization (MFO) methods. We differentiate these two genres by whether they use the optical computing system emulator models during the parameter updating process. A tabular comparison of the previous and our methods is in Tab. I. Details of the illustrations and comparisons are below.
@@
MBO methods. Physics-aware training (PAT) [15], hybrid training (HBT) [16] or adaptive training (AT) [20] (H-MBO) are one type of the MBO methods that conduct forward pass in a real system and back-propagate the gradients with the aid of a physics-based simulator. Backpropagation through learned model (L-MBO) is another method that gathers the real data from the optical computing system to train a neural proxy of the optical computing system [21, 22]. With such a learned proxy rather than the physics-based model utilized in H-MBO, the backward parameter update process can reduce the bias from the model mismatch between the simulator and the real system in L-MBO. During training, one has to optimize both the task’s and proxy’s parameters to optimize the task performance and improve the fidelity of the simulator. The task performance for the MBO, especially the L-MBO, is high while back-propagating with the digital twin also burdens the training process. The L-MBO methods implemented in optical computing are referred as hardware (camera)-in-the-loop methods. These methods have recently been used in computational display [31, 32] and camera ISP [33]. They are also known as real2sim in photolithography for solving computational lithography tasks [34, 35]. Despite the computational burden introduced by the emulation process from the digital twin, another disadvantage of the MBO methods is that all the MBO approaches require images of input objects during training, adding further burden to the overall process.
@@
MFO methods. Unlike the aforementioned MBO methods, MFO methods do not require constructing a digital twin of the optical computing system; they directly use the real system’s output or intermediate measurements to update its parameters. Population-based algorithms such as genetic algorithm (GA) methods [30] use the output from the optical computing system to optimize parameters by simulating the process of natural selection. Still, the genetic algorithm’s performance scales poorly to the dimensionality of parameters [36]. Forward-forward training (FFT) is a greedy multi-layer learning procedure for mortal computation [37] and it has been applied to train optical computing systems recently [26, 27]. However, labels must be added to each input sample during the training and testing. As a result, such training strategies require masking the input, which might be unrealistic in many scenarios, and the testing process of forward-forward training is prolonged. In Trail and error, perturbed parameters that enhance performance are retained, while those that do not are discarded, reverting to the previous state [28, 29]. This method, however, does not scale well with the dimensionality of parameters. Intermediate measurement optical backpropagation (IMOB) proposed using the adjoint variable method or direct feedback alignment, etc., to calculate gradients [23, 24, 25]. However, such methods require intermediate system measurements for every layer’s output, which might be infeasible or require additional measurement setups.
@@
We propose the gradient-based MFO (G-MFO) training method using Monte Carlo gradient estimation to update the optical computing system parameters automatically. Our work shares similarities with Zhang et al., which uses the genetic algorithm for weight update [30], as our and their methods both generate a batch of sampled parameters during the training process. The difference is that our parameter updating process is more efficient owing to a score gradient estimator rather than their heuristic hill-climbing type algorithm. Moreover, our method does not need specific conditions, such as changing the system or manipulating the input object, which are more or less required in other MBO and MFO methods.
2.2 Zeroth-Order Optimization and Its Applications to Computational Optics
The zeroth-order (ZO) optimization uses finite differences from forward evaluations instead of the first-order path-wise gradient to calculate the gradient for updating the parameters of interest [38, 39]. ZO has been utilized in many areas, such as differentiating through discontinuities in computer graphics [40], optimizing contact dynamics in robotics [41, 42], and policy learning in reinforcement learning [19]. ZO optimization has been implemented with both full-point estimation [43], and stochastic two-point estimation [44]. A significant advancement in ZO is to estimate the descent direction using multiple stochastic finite differences [45] where Monte Carlo approximation is used in the process, also called evolutionary strategies [46]. Moreover, techniques such as antithetic sampling, control variates [47], or more advanced methods to reduce variance [48, 49] in the Monte Carlo approximation process are used in multiple stochastic finite differences. Our work employs Monte Carlo gradient estimation [38], which resembles multiple stochastic finite differences.
Applications of ZO Methods to Computational Optics have been relatively sparse compared to first-order methods. In the past, there have been demonstrations in computational optics tasks such as lens design [50], auto-tuning structured light for depth estimation [51], and computer-generated holography [52]. Using only forward evaluations for gradient updates makes it beneficial for optical computing-related tasks, such as what was mentioned as optical stochastic gradient descent (SGD) in the auto-tuning structured light work [51]. The reason is that forward evaluation in optical computing is inexpensive while finding the first-order gradient is computationally costly and prone to errors. Our work highlights the advantages of using zeroth-order optimization for optimizing more complex systems, such as optical computing, and the efficacy of using Monte Carlo gradient estimation for efficient optimization.
@@ 3.3 In situ G-MFO
$=\frac{1}{M}\sum_{j=1}^{M}[\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(f_{sys}(x_{i},w_{j}),y_{i})]\nabla_{\theta}\log{{p(w_{j}|\theta)}},$ (8b)
$=\frac{1}{M}\sum_{j=1}^{M}r(w_{j})\nabla_{\theta}\log{{p(w_{j}|\theta)}}.$ (8c)
Here we define $r(w_{j})=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(f_{sys}(x_{i},w_{j}),y_{i})$ as the negative reward corresponding to each weight $w_{j}$. We use Equation 8c as the score gradient estimator for G-MFO, where the score function is $\nabla_{\theta}\log{p(w_{j}|\theta)}$, which has been widely used in other areas, such as policy gradient algorithms in reinforcement learning [19] and diffusion models [62].
@@
Figure 3: A visual illustration of the training process for the G-MFO experiment on the real system.
3.4 Experiment and Simulation Details of Our Diffractive Optical Computing System
We validate the effectiveness of the proposed G-MFO-based training strategy on a single-layer optical computing system and a two-layer diffractive optical computing system. Our work focuses on the training strategy rather than the design or adavantage of a specific optical computing system. Therefore, we do not consider nonlinearity in constructing our optical computing system, which has recently been implemented using saturation effects [20, 59, 64]. The validity of using a multi-layer versus a single-layer system without nonlinearity has been analyzed empirically in [65], where the authors attribute the improvement to increased energy resulting from the multi-layer settings. Details on the experimental setups ($f_{sys-1-layer}$, $f_{sys-2-layer}$) and the corresponding simulators ($\hat{f}_{sys-1-layer}$, $\hat{f}_{sys-2-layer}$) are in the Supplement Sec. S.1.
@@ 3.4 Experiment and Simulation Details of Our Diffractive Optical Computing System
Figure 4: Visualization of the single-layer optical computing system’s experimental outputs and confusion matrices trained with G-MFO. (a) An input phase object digit ’$2$’ from the MNIST dataset is modulated by the optical computing layer with weight $w$ trained using G-MFO. The system correctly predicts the input as digit ‘$2$’, as the output image has the largest intensity at the region corresponding to digit ’$2$’. (b) Confusion matrix on the MNIST dataset with a training accuracy of $83.1\%$. (c) An example of a ’pullover’ from the FMNIST dataset is correctly predicted. (d) Confusion matrix on the FMNIST dataset with a training accuracy of $74.0\%$.
@@ 4 Results
  • HBT: The hybrid training strategies with a physics-based simulator.
  • L-MBO: We pre-train a proxy using the task dataset to emulate the real system and then optimize the optical computing task based on the pre-trained proxy.
  • Zeroth-order optimization baselines: Including two-point estimation on a random direction [44] and full-point estimation [43, 51].
The training details for simulations and experiments are in Supplement Sec. S.3. Supplement Sec. S.6 provides a detailed description of our implementation of the HBT baseline. Supplement Sec. S.7 discusses and compares L-MBO.
@@ 4.1 General performance evaluation on the MNIST and FMNIST dataset
4.1.1 Simulation comparison between G-MFO and other zeroth-order optimization methods in simulation on a small dataset
We use a single-layer optical computing simulator to compare our method with two zeroth-order (ZO) baselines: the full-point ZO method [51, 43] and the two-point ZO method [44]. This comparison mainly evaluates the training performance, including accuracy, execution, and convergence speed. Therefore, we use a small dataset of 100 training, 100 validation, and 100 testing samples.
We present the simulation results and training time per epoch in Table II, and the curves of training accuracy in Fig. 5. The results show that sufficient samples are necessary for successful optimization. Specifically, the two-point method, with two samples per epoch, only achieves a training accuracy of $22\%$, whereas both the full-point method, using $128\times 128$ samples per epoch, and the G-MFO method, using $128$ samples per epoch, achieve a training accuracy of $99\%$. Additionally, G-MFO has improved sample efficiency than the full-point optimization method, requiring only $4\%$ of the samples to achieve an accuracy of $99\%$ compared to the full-point optimization method.
Figure 5: G-MFO balances sample efficiency and task performance for the optimization process. We compare the training curves of the two-point zeroth-order (ZO) method (green), the full-point ZO method (blue), and our G-MFO (orange) method. The full-point ZO method and G-MFO achieve $99\%$ training accuracy on the small dataset. Notably, the total number of samples of G-MFO is only approximately $4\%$ of that required by the full-point ZO method to achieve an accuracy of $99\%$. Additionally, the two-point ZO method fails to optimize effectively. The right panel presents an enlarged view of the selected region from the left panel.
Accuracy # of samples per epoch Training time per epoch
Train Val Test
Full-point [51, 43] $99.0\%$ $75.0\%$ $59.0\%$ $128\times 128$ $113~{}min$
Two-point [44] $22.0\%$ $27.0\%$ $26.0\%$ 2 $2~{}s$
G-MFO (Ours) $99.0\%$ $73.0\%$ $60.0\%$ $128$ $2~{}min$
TABLE II: Numerical comparison of G-MFO and other zeroth-order (ZO) optimization methods on FMNIST dataset on a simulated single-layer diffractive optical computing system. We show results from the full-point ZO method, the two-point ZO method, and G-MFO.
@@
TABLE III: Numerical performance comparison on MNIST and FMNIST datasets on two-layer diffractive optical computing system.
The simulation outcomes for the MNIST and FMNIST datasets are in Tab. III. The ideal testing accuracy reaches $89.5\%$ on MNIST and $80.5\%$ on FMNIST. However, the presence of misalignment reduces the accuracy to $78.4\%$ and $59.8\%$ using HBT, a decrease of approximately $10\%$ and $20\%$. In contrast, the G-MFO method hits an accuracy of $87.0\%$ and $74.1\%$, effectively mitigating the detrimental impact of system misalignment.
@@
TABLE IV: Experimental performance comparison on MNIST and FMNIST datasets on the single-layer optical computing system. Results of the ideal mode are from the simulator, whose parameters are determined from experiments while we impose no misalignment on the simulator. The lower four rows are experimental results from SBT, HBT, G-MFO, and HBT+G-MFO (G-MFO fine-tuning upon the result from HBT). Our method outperforms the SBT and HBT methods on the MNIST and FMNIST datasets.
Tab. IV quantitatively shows that our method achieves higher classification accuracy than the HBT and SBT methods on both datasets in the experiments. The SBT method performs poorly due to the reality gap between the simulator and the real system. The HBT method suffers from the bias between $\hat{f}_{sys}$ and ${f}_{sys}$ in the backward process, while the G-MFO bypasses the bias-sensitive modeling and updates gradients solely with $f_{sys}$. We further show G-MFO’s capability to fine-tune the HBT result. The last row of Tab. IV shows that the unbiased G-MFO method further improves the results of HBT. Moreover, we also empirically find that fine-tuning outperforms G-MFO only. Fig. 4 visualizes some experimental outputs and confusion matrices using the G-MFO method.
@@
We also perform experiments on training a real two-layer optical computing system $f_{sys-2-layer}$ on the MNIST dataset. For this experiment, we utilize a dataset comprising 10,000 instances for training, another 10,000 for validation, and a separate set of 10,000 for testing. The training, validation, and testing accuracy are $80.43\%$, $81.1\%$, and $80.3\%$, respectively.
Compared with the simulated G-MFO outcomes presented in Tab. III, the experimental accuracy is lower. We hypothesize that this discrepancy is attributed, in part, to mechanical disturbances encountered during the multi-day span of the experiment. Moreover, the training accuracy from a two-layer system is lower than that of a single-layer system, while the validation and testing accuracy is higher. This is because we use more samples when training a two-layer system, alleviating overfitting.
@@
4.2 Advantage: memory- and computation-efficient training enabled by G-MFO